# Optimal Coupled Sensor Placement and Path-Planning in Unknown Time -Varying Environments

### v1

@@ Abstract
We address path-planning for a mobile agent to navigate in an unknown environment with minimum exposure to a spatially and temporally varying threat field. The threat field is estimated using pointwise noisy measurements from a mobile sensor network. For this problem, we present a new information gain measure for optimal sensor placement that quantifies reduction in uncertainty in the path cost rather than the environment state. This measure, which we call the context-relevant mutual information (CRMI), couples the sensor placement and path-planning problem. We propose an iterative coupled sensor configuration and path-planning (CSCP) algorithm. At each iteration, this algorithm places sensors to maximize CRMI, updates the threat estimate using new measurements, and recalculates the path with minimum expected exposure to the threat. The iterations converge when the path cost variance, which is an indicator of risk, reduces below a desired threshold. We show that CRMI is submodular, and therefore, greedy optimization provides near-optimal sensor placements while maintaining computational efficiency of the CSCP algorithm. Distance-based sensor reconfiguration costs are introduced in a modified CRMI measure, which we also show to be submodular. Through numerical simulations, we demonstrate that the principal advantage of this algorithm is that near-optimal low-variance paths are achieved using far fewer sensor measurements as compared to a standard sensor placement method.
@@ 1 Introduction
In practical applications, however, sensor networks may be constrained by size as well as energy usage. Consider, for example, a sensor network of unmanned aerial vehicles (UAVs) for surveillance of a threat field like wildfire in a large area. Due to cost and battery limitations, it may not be possible to achieve full area coverage quickly enough to inform the actions of a ground robot to safely navigate the environment. This situation exemplifies the broader problem of path-planning with a minimal number of sensor measurements, and in turn, highlights the need for optimal sensor configuration in the context of path-planning. This problem lies at the intersection of several research areas including estimation, path-planning, and sensor placement, which we briefly review next. We note that the problem of interest here is quite different from the simultaneous localization and mapping (SLAM) problem, where the path-planning and sensing is coupled due to the assumption of fully onboard sensing. By contrast, we consider a distributed sensor network separate from the mobile agent.
Of these areas, perhaps estimation is the most mature [1]. The literature on estimation involves different techniques including Kalman filter, maximum likelihood estimator [1], and Bayesian filter [2]. Application of the extended Kalman filter (EKF), the unscented Kalman filter (UKF) [3], or the particle filter [4] is common for nonlinear dynamical systems.
@@
Path- and motion-planning are similarly mature areas of research. Generally, path-planning under uncertainty involves finding paths that minimize the expected cost. Classical approaches to path-planning include cell decomposition, probabilistic roadmaps, and artificial potential field techniques [5, 6]. Dijkstra’s algorithm, $\textrm{A}^{\ast}$, and their variants are branch-and-bound optimization algorithms that leverage heuristics to effectively steer the path search towards the goal. While classical path-planning methods are powerful, they are inherently limited by the accuracy of the environment’s available information. An accurate representation of the environment is difficult if the environment’s states or dynamics are unknown. Modern approaches to path planning leverage advanced methodologies such as adaptive informative path-planning [7], coverage path-planning [8] and informed sampling-based path planning [9]. More recently, learning based techniques, particularly deep reinforcement learning [10, 11, 12] and fuzzy logic [13] are reported for addressing environmental uncertainty.
@@
Different sensor placement approaches have been employed depending on the type of application and parameters that need to be measured. Greedy approaches based on information-based metrics [14, 15, 16] are studied. Machine learning-based sensor placement techniques are reported for efficient sensing with a minimal number of sensors and measurements as possible [17, 18]. Information theory-based sensor placement techniques utilize performance metrics that include Fisher information matrix (FIM) [19], entropy [20], Kullback-Leibler (KL) divergence [21], mutual information [22], and frame potential [14], for maximizing the amount of valuable information gathered from the surrounding environment. Similarly, [23] utilizes two information measures, one associated with mutual information based on objection detection, and another with mutual information based on classification of the detected objects. With all these performance metrics, the intention is to maximally reduce some quantification of the uncertainty. More recently, sensing and path-planning method based on reinforcement learning [24] that evaluates the performance of a called Proximal Policy Optimization (PPO) is reported. In robotics and aerospace applications, active sensing plays a crucial role in linking perception and action, enabling systems to gather meaningful information that supports intelligent decision-making. Some examples of active sensing include information-driven or cooperative active sensing [25, 26, 27], uncertainty-aware active sensing [28], and active sensing using machine learning [29, 30].
@@
In this paper we consider the problem of optimal sensor placement coupled with path-planning in an unknown dynamic environment. Specifically, we are interested in sensor placement to collect information of most relevance to the path-planning problem. The objective is to find a near-optimal path with high confidence, i.e., low variance in the path cost, with a minimal number of sensor measurements. We aim to compare such a coupled sensor configuration and path-planning (CSCP) method against decoupled methods, where sensor configuration is achieved by optimizing a metric that does not consider in any way the path-planning problem. This is a relatively new research problem. Prior works by the second author and co-workers address this problem for static (i.e., time-invariant) environments. A heuristic task-driven sensor placement approach called the interactive planning and sensing (IPAS) for static environments is reported in [34]. The IPAS method outperforms several decoupled sensor placement methods in terms of the total number of measurements needed to achieve near-optimal paths. Sensor configuration for location and field-of-view is reported in [35], also for static fields. Sensor placement for multi-agent path-planning based on entropy reduction is presented in [36].
@@ 2 Problem Formulation
Let $,{\rm Nature}$ denote the sets of real and natural numbers, respectively. We denote by $\{N\}$ the set {1, 2, …, $N$}, and by $\boldsymbol{I}_{N}$ the identity matrix of size $N$, for any $N\in{\rm Nature}.$
Consider a closed square region denoted by $\mathcal{W}\subset[2]$ and referred as the workspace, within which the mobile agent (called the actor) and a network of mobile sensors operate. In this workspace, consider a grid consisting of $N_{\mathrm{G}}$ uniformly spaced points. The coordinates of these points in a prespecified Cartesian coordinate axis system are denoted by $\boldsymbol{x}_{i}$, for each $i\in{N_{\mathrm{G}}}$. The distance between the adjacent grid points is denoted by $\delta$. The mobile agent traverses grid points according to the “4-way adjacency rule", such that the adjacent points are top, down, left, and right. Furthermore, we assume a constant speed such that the actor’s transitions to adjacent grid points occurs in a constant time step $\Delta t.$
@@
The cost $J(\boldsymbol{v})\in_{>0}$ indicates the total threat exposure for an actor on its traversal along a path $\boldsymbol{v}$ and is defined as the sum of edge transition costs, $J(\boldsymbol{v})=\sum_{\ell=1}^{L}g((v_{\ell-1},v_{\ell}),\ell\Delta t,\delta)$. The main problem of interest is to find a path $\boldsymbol{v}^{*}$ with minimum cost. Because the threat field is unknown and time-varying, its estimation is essential. A network of $N_{\mathrm{S}}$ sensors, where $N_{\mathrm{S}}\ll N_{\mathrm{G}},$ can be used to measure the intensity of threat. These sensor measurements are denoted $\boldsymbol{z}(\boldsymbol{x},t;\boldsymbol{q})=\{z_{1}(\boldsymbol{x},t;\boldsymbol{q}),z_{2}(\boldsymbol{x},t;\boldsymbol{q}),\ldots,z_{N_{s}}(\boldsymbol{x},t;\boldsymbol{q})\}.$ Sensors are placed at distinct grid points, and the set of these grid points is called the sensor configuration, $\boldsymbol{q}=\{q_{1},q_{2},\ldots,q_{N_{s}}\}\subset\{N_{\mathrm{G}}\}$.
@@
The threat field is considered to be a stochastic quantity with a predictive model involving uncertainty. Specifically, we consider a finite parametrization $c(\boldsymbol{x},t):=1+\sum_{n=1}^{N_{\mathrm{P}}}\theta_{n}(t)\phi_{n}(\boldsymbol{x})=1+\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x})\boldsymbol{\Theta}(t)$, with $\boldsymbol{\Phi}(\boldsymbol{x}):=[\phi_{1}(\boldsymbol{x})\ldots\phi_{{N_{\mathrm{P}}}}(\boldsymbol{x})]^{\intercal}$, and $\phi_{n}(\boldsymbol{x}):=\exp(-(\boldsymbol{x}-\overline{\boldsymbol{x}}_{n})^{\intercal}(\boldsymbol{x}-\overline{\boldsymbol{x}}_{n})/2a_{n})$ representing the basis functions for each $n\in\{N_{\mathrm{P}}\}$. Here, $N_{\mathrm{P}}$ is the number of parameters – i.e., the threat state, and $a_{n}$ and $\overline{\boldsymbol{x}}_{n}$ are constants. The parameter $\boldsymbol{\Theta}(t):=[\theta_{1}(t)\ldots\theta_{{N_{\mathrm{P}}}}(t)]^{\intercal}$ is to be estimated. The region $\mathcal{R}_{n}^{\mathrm{sup}}:=\left\{\boldsymbol{x}:\|\boldsymbol{x}-\boldsymbol{x}_{n}\|\leqslant 3\sqrt{a_{n}}\right\}\cap\mathcal{W}$ is defined as the region of significant support for the basis function $\phi_{n}$. Although the functions $\phi_{n}$ do not have compact support in $[2]$, their values drop below $2.5\times 10^{-3}$ outside $\mathcal{R}_{n}^{\mathrm{sup}}.$ The values of the constants $a_{n}\in_{>0}$ and $\overline{\boldsymbol{x}}_{n}\in\mathcal{W}$ are prespecified and chosen in such a manner that the combined interiors of the significant support regions cover $\mathcal{W}.$ The parameter $a_{n}$ is chosen to minimize the overlap between the basis functions, ensuring better distinction and independence among them [34].
@@
The temporal evolution of the threat is modeled by $\dot{\boldsymbol{\Theta}}(t)=A_{c}\boldsymbol{\Theta}(t)+\boldsymbol{\omega}(t)$, where $\boldsymbol{\omega}(t)\sim\mathcal{N}(0,Q_{c})$ is white process noise with $Q_{c}:=\sigma_{P}^{2}\boldsymbol{I}_{N_{\mathrm{P}}}$. Such a model may be available either from an underlying physical model of the threat, or it may be derived from data, or a combination of both. As an illustrative example, the solution to a heat diffusion equation, $\frac{\partial{c}}{\partial{t}}=\alpha(\frac{\partial^{2}{c}}{\partial{x^{2}}}+\frac{\partial^{2}{c}}{\partial{y^{2}}})$ can be approximated by $c(\boldsymbol{x},t)=1+\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x})\boldsymbol{\Theta}(t)$ such that $\boldsymbol{\Theta}(t)$ satisfies $\dot{\boldsymbol{\Theta}}(t)=\alpha\frac{\boldsymbol{\Phi}^{\intercal}}{|\boldsymbol{\Phi}|^{2}}\nabla^{2}\boldsymbol{\Phi}\boldsymbol{\Theta}(t)$, where $A_{c}:=\alpha\frac{\boldsymbol{\Phi}^{\intercal}}{|\boldsymbol{\Phi}|^{2}}\nabla^{2}\boldsymbol{\Phi}$. Discretization in time of this model is easily accomplished by the series expansion $A:=\boldsymbol{I}_{N_{\mathrm{P}}}+A_{c}\Delta t+\frac{(A_{c})^{2}(\Delta t)^{2}}{2!}+\ldots$ and $Q:=Q_{c}\Delta t+\frac{(A_{c}Q_{c}+Q_{c}A_{c}^{\intercal})(\Delta t)^{2}}{2!}+\ldots$ terminated at a desired order [1]. The discretized system dynamics are:
@@
where $\mbox{where \quad}C_{k}(\boldsymbol{q})$ $:=\left[\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{q}_{k,1}})\quad\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{q}_{k,2}})\quad\ldots\quad\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{q}_{k,N_{\mathrm{S}}}})\right]^{\intercal},$ (4)
@@
We generate stochastic estimates of the threat state with mean value $\widehat{\boldsymbol{\Theta}}(t)$ and estimation error covariance $P.$ For any path, $\boldsymbol{v}=\{v_{0},v_{1},\ldots,v_{L}\}$ in $\mathcal{G}$, the cost of the path is
$J(\boldsymbol{v}):=L+\delta\sum_{\ell=1}^{L}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\ell})\boldsymbol{\Theta}_{\ell}.$ (5)
Here, $\boldsymbol{\Theta}_{\ell}$ represent the threats at different time steps equivalent to the time steps required for an actor to follow the path $\boldsymbol{v}$. The cost $J$ becomes a random variable with distribution dependent on $\boldsymbol{\Theta}.$ Let $\mathcal{V}\subset\mathcal{W}$ be the set of grid point locations in the workspace associated with each $v_{\ell}$ in the path $\boldsymbol{v}$. The number of threat parameters $\theta_{n}(t)$, or equivalently the associated basis functions $\phi_{n}$, that lie within the path is relevant for computing the path cost $J(\boldsymbol{v})$. The set $\mathcal{K}:=\left\{m\in\{N_{\mathrm{P}}\}:\mathcal{V}\cap\mathcal{R}_{m}^{\mathrm{sup}}\neq\emptyset\right\}$ is defined as the path-relevant set of threat states for any path $\boldsymbol{v}$. In other words, $\mathcal{K}$ consists of the set of indices $m$, representing the threat parameters $\theta_{m}$ or their corresponding basis functions $\phi_{m}$, such that the path lies within the total region of significant support $\mathcal{R}_{m}^{\mathrm{sup}}$ defined by $\phi_{m}$.
An important characteristic associated with the convergence of the path-planning algorithm is the risk of the path denoted by $\rho(\boldsymbol{v})$. If $J$ is Gaussian, then the risk of the path $\boldsymbol{v}$ is defined as $\rho(\boldsymbol{v}):=\widehat{J}(\boldsymbol{v})+\sqrt{\textrm{Var}[{J}(\boldsymbol{v})]}$ [39]. Since $N_{\mathrm{S}}\ll N_{\mathrm{G}}$, it is not possible to obtain good estimates with only one set of measurements, and it is required to take measurements repeatedly over a finite number of iterations.

### v2

@@ Abstract
We address path-planning for a mobile agent to navigate in an unknown environment with minimum exposure to a spatially and temporally varying threat field. The threat field is estimated using pointwise noisy measurements from a mobile sensor network. For this problem, we present a new information gain measure for optimal sensor placement that quantifies reduction in uncertainty in the path cost rather than the environment state. This measure, which we call the context-relevant mutual information (CRMI), couples the sensor placement and path-planning problem. We propose an iterative coupled sensor configuration and path-planning (CSCP) algorithm. At each iteration, this algorithm places sensors to maximize CRMI, updates the threat estimate using new measurements, and recalculates the path with minimum expected exposure to the threat. The iterations converge when the path cost variance, which is an indicator of risk, reduces below a desired threshold. We show that CRMI is submodular, and therefore, greedy optimization provides near-optimal sensor placements while maintaining computational efficiency of the CSCP algorithm. Distance-based sensor reconfiguration costs are introduced in a modified CRMI measure, which we also show to be submodular. Through numerical simulations, we demonstrate that the principal advantage of this algorithm is that near-optimal low-variance paths are achieved using far fewer sensor measurements as compared to a standard sensor placement method. footnotetext: Preliminary results from this work were presented at the 2025 AIAA SciTech Forum and Exposition, January 6 - 10, Orlando, FL, USA. AIAA paper number 2025-2069.
@@ 1 Introduction
Figure 1 provides a motivating example of a time-varying flood map based on data collected during Hurricane Harvey at spatially distributed gauge stations over 15-minute time intervals [1]. Initially, there is low water discharge with no signs of flooding in Fig. 1(a). Significant flooding is visible in Figure 1(c) at $t=1500$ minutes, which subsequently recedes. The changing flood levels may be considered as a spatiotemporally-varying “threat” to which exposure of, say, an emergency first-response vehicle, should be minimized.
@@
(a) $t=500$ min.
(b) $t=1000$ min.
(c) $t=1500$ min.
(d) $t=2000$ min.
Figure 1: Visualization of the flood map at different times (minutes).
@@
In practical applications, however, sensor networks may be constrained by size as well as energy usage. In the flood evolution example, a sensor network of unmanned aerial vehicles (UAVs) may be depoyed for real-time surveillance of flooding. Due to cost and battery limitations, it may not be possible to achieve full area coverage quickly enough to inform the actions of a ground robot to safely navigate the environment. This situation exemplifies the broader problem of path-planning with a minimal number of sensor measurements, and in turn, highlights the need for optimal sensor configuration in the context of path-planning. This problem lies at the intersection of several research areas including estimation, path-planning, and sensor placement, which we briefly review next. We note that the problem of interest here is quite different from the simultaneous localization and mapping (SLAM) problem, where the path-planning and sensing is coupled due to the assumption of fully onboard sensing. By contrast, we consider a distributed sensor network separate from the mobile agent.
Of these areas, perhaps estimation is the most mature [2]. The literature on estimation involves different techniques including Kalman filter, maximum likelihood estimator [2], and Bayesian filter [3]. Application of the extended Kalman filter (EKF), the unscented Kalman filter (UKF) [4], or the particle filter [5] is common for nonlinear dynamical systems. Several data-driven estimation techniques have been employed for spatiotemporal modeling of hazardous regions, particularly in robotics, environmental monitoring, and surveillance. These include a supervised learning approach using Gaussian process regression for threat field estimation [6], a statistical generalized additive model for landslide hazard estimation [7] and and a combined approach using Bayesian inference and random forest for the spatial prediction of wildfires [8].
@@
Path- and motion-planning are similarly mature areas of research. Generally, path-planning under uncertainty involves finding paths that minimize the expected cost. Classical approaches to path-planning include cell decomposition, probabilistic roadmaps, and artificial potential field techniques [9, 10]. Dijkstra’s algorithm, $\textrm{A}^{\ast}$, and their variants are branch-and-bound optimization algorithms that leverage heuristics to effectively steer the path search towards the goal. While classical path-planning methods are powerful, they are inherently limited by the accuracy of the environment’s available information. An accurate representation of the environment is difficult if the environment’s states or dynamics are unknown. Modern approaches to path planning leverage advanced methodologies such as adaptive informative path-planning [11], coverage path-planning [12] and informed sampling-based path planning [13]. More recently, learning based techniques, particularly deep reinforcement learning [14, 15, 16, 17] and fuzzy logic [18] are reported for addressing environmental uncertainty. In risk-aware path planning, the objective is not only to find a feasible or optimal path but also to minimize exposure to uncertain or hazardous regions [19]. Recent risk-aware path planning techniques integrate reinforcement learning with failure mode and effect analysis to ensure safe and complete coverage in hazardous environments [20].
@@
Different sensor placement approaches have been employed depending on the type of application and parameters that need to be measured. Greedy approaches based on information-based metrics are presented in [21, 22, 23]. Machine learning-based sensor placement techniques are reported for efficient sensing with a minimal number of sensors and measurements as possible [24, 25]. Information theory-based sensor placement techniques utilize performance metrics such as the Fisher information matrix (FIM) [26], entropy [27], Kullback-Leibler (KL) divergence [28], mutual information [29], and frame potential [21] to maximize the amount of valuable information gathered from the surrounding environment. Similarly, [30] utilizes two information measures, one associated with mutual information based on objection detection, and another with mutual information based on classification of the detected objects. With all these performance metrics, the intention is to maximally reduce some quantification of the uncertainty. More recently, a sensing and path-planning method based on reinforcement learning has been reported, which evaluates performance using a technique called proximal policy optimization (PPO) [31]. In robotics and aerospace applications, active sensing plays a crucial role in linking perception and action, enabling systems to gather meaningful information that supports intelligent decision-making. Some examples of active sensing include information-driven or cooperative active sensing [32, 33, 34], uncertainty-aware active sensing [35], and active sensing using machine learning [36, 37].
@@
In this paper we consider the problem of optimal sensor placement coupled with path-planning in an unknown dynamic environment. Specifically, we are interested in sensor placement to collect information of most relevance to the path-planning problem. The objective is to find a near-optimal path with high confidence, i.e., low variance in the path cost, with a minimal number of sensor measurements. We aim to compare such a coupled sensor configuration and path-planning (CSCP) method against decoupled methods, where sensor configuration is achieved by optimizing a metric that does not consider the path-planning problem in any way. This is a relatively new research problem. Prior works by the second author and co-workers address this problem for static (i.e., time-invariant) environments. A heuristic task-driven sensor placement approach called the interactive planning and sensing (IPAS) for static environments is reported in [41]. The IPAS method outperforms several decoupled sensor placement methods in terms of the total number of measurements needed to achieve near-optimal paths. Sensor configuration for location and field-of-view is reported in [6], also for static fields. Sensor placement for multi-agent path-planning based on entropy reduction is presented in [42].
@@ 2 Problem Formulation
Let $\rm Nature$} denote the sets of real and natural numbers, respectively. We denote by $\{N\}$ the set {1, 2, …, NN}, and by $\boldsymbol{I}_{N}$ the identity matrix of size $N$, for any $N\in{\rm Nature}.$
Consider a closed square region denoted by $\mathcal{W}\subset[2]$ and referred as the workspace, within which the mobile agent (called the actor) and a network of mobile sensors operate. In this workspace, consider a grid consisting of $N_{\mathrm{G}}$ uniformly spaced points. The coordinates of these points in a prespecified Cartesian coordinate axis system are denoted by $\boldsymbol{x}_{i}=(x_{i},y_{i})$, for each $i\in{N_{\mathrm{G}}}$. The distance between the adjacent grid points is denoted by $\delta$. The mobile agent traverses grid points according to the “4-way adjacency rule", such that the adjacent points are top, down, left, and right. Furthermore, we assume a constant speed such that the actor’s transitions to adjacent grid points occurs in a constant time step $\Delta t.$
@@
The cost $J(\boldsymbol{v})\in_{>0}$ indicates the total threat exposure for an actor on its traversal along a path $\boldsymbol{v}$ and is defined as the sum of edge transition costs, $J(\boldsymbol{v})=\delta\sum_{\ell=1}^{L}g((v_{\ell-1},v_{\ell}),\ell\Delta t)$. The main problem of interest is to find a path $\boldsymbol{v}^{*}$ with minimum cost. Because the threat field is unknown and time-varying, its estimation is essential. A network of $N_{\mathrm{S}}$ sensors, where $N_{\mathrm{S}}\ll N_{\mathrm{G}},$ can be used to measure the intensity of threat. These sensor measurements are denoted $\boldsymbol{z}(\boldsymbol{x},t;\boldsymbol{q})=\{z_{1}(\boldsymbol{x},t;q_{1}),z_{2}(\boldsymbol{x},t;q_{2}),\ldots,z_{N_{s}}(\boldsymbol{x},t;q_{N_{s}})\}.$ Sensors are placed at distinct grid points, and the set of these grid points is called the sensor configuration, $\boldsymbol{q}=\{q_{1},q_{2},\ldots,q_{N_{s}}\}\subset\{N_{\mathrm{G}}\}$.
@@
The threat field is considered to be a stochastic quantity with a predictive model involving uncertainty. Specifically, we consider a finite parametrization $c(\boldsymbol{x},t):=1+\sum_{n=1}^{N_{\mathrm{P}}}\theta_{n}(t)\phi_{n}(\boldsymbol{x})=1+\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x})\boldsymbol{\Theta}(t)$, with $\boldsymbol{\Phi}(\boldsymbol{x}):=[\phi_{1}(\boldsymbol{x})\ldots\phi_{{N_{\mathrm{P}}}}(\boldsymbol{x})]^{\intercal}$, and $\phi_{n}(\boldsymbol{x}):=\exp(-(\boldsymbol{x}-\overline{\boldsymbol{x}}_{n})^{\intercal}(\boldsymbol{x}-\overline{\boldsymbol{x}}_{n})/2a_{n})$ representing the basis functions for each $n\in\{N_{\mathrm{P}}\}$. Here, $N_{\mathrm{P}}$ denotes the number of parameters (or bases), representing the threat state, and $a_{n}$ and $\overline{\boldsymbol{x}}_{n}$ are constants. The locations of the basis functions $\boldsymbol{\Phi}$ are fixed and remain unchanged throughout the entire sensing and path planning process. Although the actor has prior knowledge of the functional forms of these functions, the threat parameter $\boldsymbol{\Theta}(t):=[\theta_{1}(t)\ldots\theta_{{N_{\mathrm{P}}}}(t)]^{\intercal}$ is unknown and must be estimated. The values of the constants $a_{n}\in_{>0}$ and $\overline{\boldsymbol{x}}_{n}\in\mathcal{W}$ are prespecified and chosen in such a manner that the combined interiors of the significant support regions cover $\mathcal{W}.$ The parameter $a_{n}$ is chosen to minimize the overlap between the basis functions, ensuring better distinction and independence among them [41]. In general, the basis functions should be chosen to approximate threat field data, e.g., the flood map data in Fig. 1.
@@
The temporal evolution of the threat is modeled by $\dot{\boldsymbol{\Theta}}(t)=A_{c}\boldsymbol{\Theta}(t)+\boldsymbol{\omega}(t)$, where $\boldsymbol{\omega}(t)\sim\mathcal{N}(0,Q_{c})$ is white process noise with $Q_{c}:=\sigma_{P}^{2}\boldsymbol{I}_{N_{\mathrm{P}}}$. Such a model may be available either from an underlying physical model of the threat, or it may be derived from data, or a combination of both. As an illustrative example, the solution to a heat diffusion equation, $\frac{\partial{c}}{\partial{t}}=\alpha(\frac{\partial^{2}{c}}{\partial{x^{2}}}+\frac{\partial^{2}{c}}{\partial{y^{2}}})$ can be approximated by $c(\boldsymbol{x},t)=1+\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x})\boldsymbol{\Theta}(t)$ such that $\boldsymbol{\Theta}(t)$ satisfies $\dot{\boldsymbol{\Theta}}(t)=\alpha\frac{\boldsymbol{\Phi}^{\intercal}}{|\boldsymbol{\Phi}|^{2}}\nabla^{2}\boldsymbol{\Phi}\boldsymbol{\Theta}(t)$, where $A_{c}:=\alpha\frac{\boldsymbol{\Phi}^{\intercal}}{|\boldsymbol{\Phi}|^{2}}\nabla^{2}\boldsymbol{\Phi}$.
We restrict the scope of this paper to linear threat field dynamics for the purpose of establishing a proof of convergence of the proposed method. However, the method itself is not limited to linear dynamics. We implement the Unscented Kalman Filter in our current threat field estimator in anticipation of this future extension. Our current CSCP implementation will work as-is for nonlinear threat dynamics, but the CRMI calculations will be approximations. Other works in the literature, e.g., [45] have studied MI calculations for nonlinear systems, which we can easily adapt to CRMI in the future.
Discretization in time of this model is easily accomplished by the series expansion $A:=\boldsymbol{I}_{N_{\mathrm{P}}}+A_{c}\Delta t+\frac{(A_{c})^{2}(\Delta t)^{2}}{2!}+\ldots$ and $Q:=Q_{c}\Delta t+\frac{(A_{c}Q_{c}+Q_{c}A_{c}^{\intercal})(\Delta t)^{2}}{2!}+\ldots$ terminated at a desired order [2]. The discretized system dynamics are:
@@
$\mbox{where \quad}C_{k}(\boldsymbol{q})$ $:=\left[\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{q}_{k,1}})\quad\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{q}_{k,2}})\quad\ldots\quad\boldsymbol{\Phi}(\boldsymbol{x}_{\boldsymbol{q}_{k,N_{\mathrm{S}}}})\right]^{\intercal},$ (4)
@@
We generate stochastic estimates of the threat state with mean value $\widehat{\boldsymbol{\Theta}}_{k}$ and estimation error covariance $P.$ For any path, $\boldsymbol{v}=\{v_{0},v_{1},\ldots,v_{L}\}$ in $\mathcal{G}$, the cost of the path is
$J(\boldsymbol{v}):=\delta{\sum_{\ell=1}^{L}}c(\boldsymbol{x}_{\boldsymbol{v}_{\ell}},t)=\delta\left(L+{\sum_{\ell=1}^{L}}\boldsymbol{\Phi}^{\intercal}(\boldsymbol{x}_{\boldsymbol{v}_{\ell}})\boldsymbol{\Theta}_{\ell}\right).$ (5)
Here, $\boldsymbol{\Theta}_{\ell}$ represent the threats at different iterations (or time steps) equivalent to the iterations required for an actor to follow the path $\boldsymbol{v}$. The cost $J$ becomes a random variable with distribution dependent on $\boldsymbol{\Theta}.$ If $\boldsymbol{\Theta}$ is Gaussian, then $J$ is also Gaussian because it is linearly dependent on $\boldsymbol{\Theta}.$ Let $\mathcal{V}\subset\mathcal{W}$ be the set of grid point locations in the workspace associated with each $v_{\ell}$ in the path $\boldsymbol{v}$. Note that while $\theta_{n}(t)$ and $\phi_{n}$ represent the total number of threat parameters and the basis functions involved in the threat field generation, $\theta_{m}(t)$ and $\phi_{m}$ denote the subset of threat parameters and associated basis functions that lie within the path and are relevant for computing the path cost $J(\boldsymbol{v})$. The set $\mathcal{K}:=\left\{m\in\{N_{\mathrm{P}}\}:\mathcal{V}\cap\mathcal{R}_{m}^{\mathrm{sup}}\neq\emptyset\right\}$ is defined as the path-relevant set of threat states for any path $\boldsymbol{v}$. In other words, $\mathcal{K}$ consists of the set of indices $m$, representing the threat parameters $\theta_{m}$ or their corresponding basis functions $\phi_{m}$, such that the path lies within the total region of significant support $\mathcal{R}_{m}^{\mathrm{sup}}$ defined by $\phi_{m}$.
An important characteristic associated with the convergence of the path-planning algorithm is the risk of the path denoted by $\rho(\boldsymbol{v})$. For a Gaussian $J$, the risk of the path $\boldsymbol{v}$ is defined as $\rho(\boldsymbol{v}):=\widehat{J}(\boldsymbol{v})+\sqrt{\textrm{Var}[{J}(\boldsymbol{v})]}$ [46]. Since $N_{\mathrm{S}}\ll N_{\mathrm{G}}$, it is not possible to obtain good estimates with only one set of measurements, and it is required to take measurements repeatedly over a finite number of iterations.